OpenStudy (anonymous):

Furnace oil has a heat of combustion of 44 MJ/kg. assuming that 70% of the heat is useful, how many kilograms of oil are required to raise the temperature of 2000 kg of water from 20 degree Celsius to 99 degree celsius

OpenStudy (fifciol):

\[0.7 *44*10^6=\frac{C_w\Delta T m_w}{m_o}\] \[m_o=\frac{ 4189,9*79*2000 }{0,7*44*10^6 }\]

OpenStudy (fifciol):

combustion is the energy per mass of oil. We know that 70% of that is useful so we multiply it by 0,7. The heat is C_w * mass of water * increase of temparature, and we have to divide that by mass of an oil to get the same unit as on left side.

OpenStudy (fifciol):

70%=0,7

OpenStudy (anonymous):

how did you convert MJ to J?

OpenStudy (fifciol):

MJ=10^6 J

OpenStudy (anonymous):

oh.. can you please explain how did you come up with the formula..

OpenStudy (fifciol):

combustion=heat produced \ mass of oil

OpenStudy (anonymous):

thanks. i get it now :)

OpenStudy (fifciol):

yw